### v1

# Evaluating Perspectival Biases in Cross-Modal Retrieval

Abstract
Multimodal retrieval systems are expected to operate in a semantic space, agnostic to the language or cultural origin of the query. In practice, however, retrieval outcomes systematically reflect perspectival biases: deviations shaped by linguistic prevalence and cultural associations. We study two such biases. First, prevalence bias refers to the tendency to favor entries from prevalent languages over semantically faithful entries in image-to-text retrieval. Second, association bias refers to the tendency to favor images culturally associated with the query over semantically correct ones in text-to-image retrieval. Results show that explicit alignment is a more effective strategy for mitigating prevalence bias. However, association bias remains a distinct and more challenging problem. These findings suggest that achieving truly equitable multimodal systems requires targeted strategies beyond simple data scaling and that bias arising from cultural association may be treated as a more challenging problem than one arising from linguistic prevalence.
Keywords model bias/fairness evaluation $\cdot$ multimodality $\cdot$ multilingual evaluation $\cdot$ language/cultural bias analysis
1 Introduction
As Nietzsche [1] observed, “there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere. Large models inherit this perspectival character through their training data; what they learn to represent depends on the frequency of appearance and co-occurrence structure. As a result, the latent space of such models does not always function as the robust, language-agnostic semantic space we expect. Instead, retrieval outcomes can be skewed, favoring linguistic prevalence or cultural association over true semantic relevance. The effect of such a perspectival character on both image-to-text and text-to-image retrievals is illustrated in Figure 1. Understanding and quantifying these effects is crucial for ensuring consistent retrieval performance across languages and cultures.
Figure 1: Two Forms of Perspectival Biases. (a) Prevalence bias: an image query favors high-resource languages. A retrieval model places English results above semantically equivalent Japanese and Thai captions. (b) Association Bias: A visualized model’s embedding space, demonstrating how a Japanese text query for “necklace” retrieves culturally proximate images (Japanese masks) instead of the semantically correct one (Kenyan necklace).
Multimodal retrieval enables cross-modality search, primarily between text and images. Early models, such as CLIP [2], align vision and language representations through paired supervision. Recent Multimodal Large Language Models (MLLMs) [3, 4] achieve alignment implicitly through large-scale pretraining. Despite these advancements, the critical issue of language and cultural bias in retrieval remains underexplored.
This lack of study is concerning given that state-of-the-art retrievers are trained on web-scale, text-image datasets like LAION [5] and WebLi [6], which are overwhelmingly English-centric. While these datasets are constructed using English alt-text, images with high cultural specificity often retain alt-text in their native languages. As observed in multilingual food datasets [7], items like the Catalan pastry “coca de recapte” are exclusively described in Catalan or Spanish. This might allow models to develop emergent multilingual capabilities, but it also risks introducing a systemic bias where the model learns spurious correlations, preferentially matching images with text from a specific majority or “expected” language.
A key barrier to investigating these biases is the absence of targeted metrics and benchmarks designed to quantify them. To address this gap, we introduce an evaluation framework for both retrieval directions, each capturing a distinct form of perspectival bias. Image-to-text retrieval. In the absence of linguistic cues, retrievals reveal how the prevalence of certain languages in the training data shapes the results. To assess this prevalence bias, we propose the Discounted Language Bias Kullback–Leibler Divergence (DLBKL), inspired by Language Bias Kullback–Leibler Divergence (LBKL) [8], which measures how strongly retrieval relevance depends on language rather than semantics, as shown in Figure 1a. Text-to-image retrieval. When linguistic and cultural cues are present in the query, retrievals reveal the model’s tendency to favor culturally associated visual patterns over semantically aligned ones. We term this association bias and construct a balanced, cross-cultural, and cross-lingual dataset to disentangle semantic relevance from cultural proximity, as shown in Figure 1b.
Using these tools, we conduct an empirical analysis comparing the perspectival biases inherent in retrievers adapted from MLLMs with those trained using explicit cross-lingual alignment techniques [9, 10]. Our findings reveal that models with explicit alignment mechanisms exhibit lower biases, highlighting a critical trade-off between the scale of MLLMs and the fairness of more targeted alignment strategies.
We summarize our contributions as follows: (i) We propose DLBKL, a metric for quantifying prevalence bias in multimodal retrieval within a multilingual candidate pool to assess language fairness. (ii) We introduce a novel vision-language dataset, parallel across culture and language, designed to assess association bias.
2 Related work
2.1 Multimodal Retrievers
Retrieving image information using a text query can be accomplished by two methods: sparse and dense retrieval. Sparse retrieval methods utilize a high-dimensional representation extracted from words in a text query or an image caption [11, 12]. While these methods are fast, they do not understand the semantics of the image as they solely rely on the image caption to represent the content. To handle the challenge of understanding semantic meaning, deep learning-based dense retrieval methods have been developed. For example, CLIP [2] and ALIGN [13] used a dual-encoder architecture specifically trained to connect the semantic meanings of text-image pairs using contrastive loss objectives. Both models were built on similar principles but varied in text-image data size and utilized different text/image encoders. The recent ColQwen [14] and GME [15] model adapts a Large Language Model (LLM) to learn a more intensive semantic connection between text and images, converting them to a multimodal retrieval model.
2.2 Language Bias in Multi-model Retrievers
Language bias in multimodal retrieval refers to performance differences that arise when semantically equivalent queries in different languages yield divergent rankings, often favoring high-resource languages such as English. Prior work frames this along two fairness axes: (i) an individual-level notion, where multilingual variants of the same query should produce similar results, and (ii) a group-level notion, where aggregate retrieval performance should remain balanced across languages [16].
A study on multilingual retrieval benchmark [17] reports uneven performance across languages, with comparatively stronger results on English and other high-resource languages for various modern multimodal retrievers, despite their large scale in data and parameters. For instance, some models exhibit significant variation in NDCG [18] scores across different languages, indicating that retrieval effectiveness is not uniform.
To quantify such disparities, fairness-aware metrics from ranking literature, such as exposure parity [18], have been adapted to language as a protected attribute. More recently, Adewumi et al. [19] surveyed multimodal bias, emphasizing the lack of dedicated language-focused evaluation protocols. Addressing this, Laosaengpha et al. [8] proposed LBKL, a distributional measure of divergence between retrieval results across language variants. Although LBKL was designed for measuring text modality bias, it can technically be extended to multimodal retrieval, enabling a more fine-grained detection of scores across languages. These works highlight several metrics for measuring language bias in multimodal retrievers. However, despite their effectiveness in measuring language bias, none take retrieval rankings into account.
2.3 Multilingual and Cross-Lingual Retrieval Strategies
Two dominant paradigms exist for building multilingual multimodal retrieval systems: holistic end-to-end pre-training and explicit cross-lingual alignment.
Holistic End-to-End Pre-training on Large-Scale Multilingual Data
This approach, anchored by foundation MLLMs like Qwen2.5-VL [3], aims to learn an emergent universal representation from web-scale, mixed-language data. Within this paradigm, some models like GME [15] and the standard ColQwen series [14] are fine-tuned on predominantly English datasets. Others, such as the multilingual ColQwen series [14] and jina-embeddings-v4 [20], intentionally incorporate extensive multilingual data to improve fairness.
Explicit Cross-Lingual Alignment via Knowledge Distillation
This alternative strategy uses knowledge distillation to align text encoders for new languages to a strong, pre-existing English model’s embedding space, such as CLIP’s. This data-efficient method, exemplified by M-CLIP [10], typically requires only parallel text corpora to force non-English embeddings to mimic their English counterparts via a teacher-student setup.
Our work evaluates models representing both paradigms, providing a direct comparison of biases inherent to each approach.
3 Methodology for Evaluating Bias in Multimodal Retrieval
This section outlines the framework developed to investigate perspectival bias in multilingual, multimodal retrieval systems. We first state our guiding research questions and then explain the studies that address these research questions in Sections 3.1 and 3.2. Our investigation centers on two complementary perspectives, corresponding to different retrieval directions, as shown in Figure 2:
  1. RQ1
[…]
  2. RQ2
[…]
Figure 2: Overview of the study. First, in RQ1, we identify language prevalence in image-to-text retrieval by analyzing the language of the retrieved text and comparing it with high-resource languages, as well as medium- and low-resource languages. Second, in RQ2, we identify the association bias using self-preference behavior of the model by retrieving an image with three candidates: semantically relevant, culturally relevant, and a non-relevant candidate.
3.1 RQ1: Image-to-Text Retrieval Study
In this study, we assess the bias arising from linguistic prevalence by examining the discrepancy between an expected “fair” language distribution and the observed one. That is, the discrepancy should be zero if the linguistic prevalence has no effect on the retrieval results and increases as the results deviate from the ideal case. As discussed in Section 2.2, existing work lacks a dedicated metric to quantify such a discrepancy in multimodal retrieval.
As the first step to closing this gap, we apply the Language Bias Kullback–Leibler (LBKL) Divergence proposed by Laosaengpha et al. [8], which measures the divergence between an expected language distribution and the observed distribution in a retrieved list. Given proportions of language A and B in the ground truth ($P_{A}(x),P_{B}(x)$) and in the retrieved set ($Q_{A}(x),Q_{B}(x)$), LBKL is given as:
$\text{LBKL}=\frac{\sum\limits_{i=1}^{q}\left[P_{\text{A}}(x)\log\frac{P_{\text{A}}(x)}{Q_{\text{A}}(x)}+P_{\text{B}}(x)\log\frac{P_{\text{B}}(x)}{Q_{\text{B}}(x)}\right]}{q}$ (1)
While LBKL can be applied to cross-modal retrieval, it is rank-agnostic: deviations at rank 1 are penalized equally to deviations at rank 100. This underestimates the harm in systems that concentrate resource-driven bias near the top ranks.
For the next step, we extend LBKL by introducing the Discounted Language Bias Kullback-Leibler (DLBKL) divergence, which incorporates a logarithmic rank discount inspired by NDCG [18]. We assign a weight $w(i)=1/\log_{2}(i+1)$ to each rank $i$. The rank-weighted proportion for a language $l$ is then:
$Q^{\prime}_{l}(x)=\frac{\sum_{i=1}^{k}w(i)\cdot\mathbb{I}(\text{doc}_{i}\text{ is }l)}{\sum_{i=1}^{k}w(i)}$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function. DLBKL is calculated by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in the LBKL formula. As illustrated in Figure 3, DLBKL penalizes top-ranked disparities more heavily, aligning the metric with user exposure and better capturing the discrepancy between the ideal case and observed one in multimodal retrieval.
Figure 3: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists where high-resource languages dominate the top ranks.
3.2 RQ2: Text-to-Image Retrieval Study
Figure 4: Overview of the XCM dataset creation process, designed to produce a benchmark with parallelism across semantics, cultures, and languages.
To quantify the degree to which models prioritize cultural association over semantic fidelity, a phenomenon we illustrate in Figure 1(b), a benchmark with a parallel structure in its cultural dimension is necessary. To the best of our knowledge, no such benchmark exists, so we make two primary contributions. First, we construct and introduce the Cross-Cultural Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose. Second, we propose the Self-Preference Cultural Bias Score (SP), a new metric for explicitly measuring this form of bias.
3.2.1 The 3XCM Dataset Benchmark
To evaluate association bias, we constructed the 3XCM benchmark [^1] Research release only (CC BY-NC-SA 4.0). Ethical review required for production use.
Available at: https://huggingface.co/datasets/Chula-AI/association_bias_benchmark.
. The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
The image gathering stage, summarized in Figure 4, consisted of three steps:
  • Concept Generation We used Gemini [^2] Version used: gemini-2.5-flash (Released June 17, 2025). to generate a large pool of concepts, which we manually curated to a final set of 138 coarse-grained, culturally-inclusive concepts (e.g., "train", "food"). Each concept is an abstract, semantic category that uses shared properties to group a broad, culturally-inclusive range of entities. The prompt for generating concepts can be found in Appendix A.
  • Concept De-duplication: We use BGE-M3 [21] to de-duplicate concepts based on similarity with a threshold of 0.92.
  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [22] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
  • Image De-duplication: To ensure visual diversity, we performed two-stage de-duplication within each concept. First, near-exact duplicates were removed automatically using an embedding model. Subsequently, three human annotators, following the guidelines in Appendix F, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
Leveraging the collected cultural images, we introduce a novel evaluation paradigm that employs a forced-choice task. This setup is designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association. As illustrated in Figure 5, for a given query (e.g., "food" in Thai), the model is presented with a triplet of image candidates: (i) Semantically Relevant: same concept, different culture (e.g., Nigerian food); (ii) Culturally Relevant: different concept, same culture (e.g., Thai traditional dance); and (iii) Non-Relevant: different concept and culture (e.g., Japanese gas station).
Figure 5: Illustration of association bias evaluation. A Thai text query for “food” is evaluated against three candidates designed to isolate semantic faithfulness vs. cultural relevance.
The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix J and N respectively.
3.2.2 Self-Preference Cultural Bias Score (SP)
With the constructed dataset, we can now measure the discrepancy between the ideal case and the observed one, where bias arising from cultural association may intervene. Ideally, the discrepancy should be zero when image retrieval depends solely on semantic relevance, and it should increase as the model’s preference tends towards images associated with the culture of the query, rather than semantic accuracy. To quantify the discrepancy, we propose a metric called the self-preference cultural bias score (SP), which can be computed as follows:
$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{sem},i},S_{\text{cul},i},S_{\text{non},i})\right)$ (3)
$\text{SP}=\frac{M_{\text{cul}}}{M_{\text{sem}}}$ (4)
where $M_{k}$ is the proportion of times a candidate of type $k$ receives the highest similarity score across $N$ total trials. The candidate type $k$ can be semantically relevant (sem), culturally relevant (cul), or non-relevant (non). The similarity score for candidate type $k$ in trial $i$ is denoted by $S_{k,i}$. The indicator function $\mathbb{I}(\cdot)$ is 1 if the condition is true and 0 otherwise. The SP score (Eq. 4) is then the ratio of cultural wins ($M_{\text{cul}}$) to semantic wins ($M_{\text{sem}}$). In this way, a higher SP score indicates stronger cultural self-preference over semantic faithfulness and thus a greater extent of association bias.
4 Experimental Setup
To answer our research questions, we conducted two main experiments. For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [23]. The dataset offers a balanced multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities. We evaluate models using Accuracy@5, NDCG@10, LBKL@10, and our proposed DLBKL@10. For RQ2, we performed text-to-image retrieval on our newly created XCM benchmark, evaluating models using our proposed SP score.
For both RQ1 and RQ2, we selected a representative suite of models spanning three distinct architectural paradigms, as shown in Table 1:
  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data. We include the original CLIP-L/14 as a powerful baseline, and Chinese-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space. We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
  • MLLM-Based Retrieval Embedders: This modern paradigm adapts large, pre-trained Multimodal Language Models for retrieval. We evaluate several state-of-the-art models, including the ColQwen series (v0.2, 3b-M, 7b-M), GME models (Qwen2-2B, Qwen2-7B), and Jina-E-v4.
Full model identifiers are available in Appendix H.
5 Experimental Results
Our experiments are designed to provide empirical examinations of perspectival biases manifested in image-to-text and text-to-image retrievals.
5.1 Image-to-Text Evaluation (RQ1)
All models exhibit some degree of linguistic prevalence bias. For most MLLM-based models (Jina, ColQwen, GME), the DLBKL score is higher than the LBKL score, as shown in Table 1.This confirms that bias is more pronounced at the top of the ranked list, as these models tend to rank results from medium-to-high resource languages. Results for additional ranks and an example of retrieval result can be found in Appendix D.
Model Acc LBKL DLBKL NDCG
[…]
Vision-Language Contrastive Models
CLIP-L/14 0.509 5.673 5.684 0.290
Chinese-CLIP-L/14 0.355 5.046 5.055 0.207
Cross-lingual Alignment Models
XLM-R-L/14 0.924 0.320 0.333 0.736
XLM-R-B/16plus 0.968 0.110 0.125 0.791
MLLM-Based Retrieval Embedders
ColQwen2.5-3b-M 0.894 0.792 0.817 0.605
ColQwen2.5-7b-M 0.926 0.821 0.849 0.665
ColQwen2.5-v0.2 0.754 3.834 3.867 0.481
GME-Qwen2-2B 0.967 3.121 3.174 0.717
GME-Qwen2-7B 0.979 1.371 1.420 0.770
Jina-E-v4 0.972 0.915 0.951 0.775
Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL. Explicit alignment models (XLM-R) show substantially lower bias.
This phenomenon is visualized in Figure 6, which shows a clear dominance of high-resource languages in the top ranks. This further illustrates the overall disparity in retrieval frequency between language resource tiers as shown in Figure 7.
Figure 6: Distribution of language groups across retrieval ranks. High-resource languages (blue) dominate the top ranks, a bias captured by DLBKL.
Figure 7: Histogram of retrieved language frequencies. MLLM-based models disproportionately retrieve texts from medium-high resource languages (blue) over low-resource ones (orange).
Crucially, the explicit alignment models (XLM-R series) achieve the lowest bias scores by a significant margin, with XLM-R-B/16plus demonstrating near-zero linguistic prevalence bias according to both metrics, while maintaining high retrieval accuracy. This provides strong initial evidence that direct alignment is a more effective strategy for enforcing language fairness than relying on emergent capabilities from large-scale pre-training.
Building on these observations, we note that LBKL and DLBKL quantify distributional bias rather than relevance, and therefore need not correlate with accuracy or NDCG in Table 1. To assess both correctness and fairness, these bias metrics should be interpreted jointly with accuracy (and/or NDCG). Finally, while LBKL/DLBKL capture cross-language imbalance, they do not measure model self-preference (e.g., favoring the query language over others); we operationalize and evaluate that phenomenon with our SP score.
5.2 Text-to-Image Evaluation (RQ2)
Using the proposed XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models. In this evaluation, the semantic win rate ($M_{\text{sem}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias. We observe that the baseline CLIP and Chinese-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
[…]
Vision-Language Contrastive Models
CLIP-L/14 51.24% 40.78% 7.98% 0.80
Chinese-CLIP-L/14 56.39% 31.65% 11.95% 0.56
Cross-lingual Alignment Models
XLM-R-L/14 85.53% 6.84% 7.63% 0.08
XLM-R-B/16plus 87.54% 6.23% 6.24% 0.07
LLM-Based Retrieval Embedders
GME-Qwen2-2B 83.34% 11.64% 5.02% 0.14
GME-Qwen2-7B 84.63% 11.26% 4.11% 0.13
ColQwen2.5-v0.2 82.10% 10.93% 6.97% 0.13
ColQwen2.5-3B-M 83.36% 10.65% 6.00% 0.13
ColQwen2.5-7B-M 84.07% 11.40% 4.53% 0.14
Jina-E-v4 87.56% 7.20% 5.24% 0.08
Table 2: Results on the XCM benchmark for Self-Preference Cultural Bias.
Our culture-specific analysis reveals that this self-preference is a symptom of missing linguistic knowledge, as shown in Figure 8. The CLIP-L/14 model, lacking a robust understanding of non-Latin scripts, defaults to matching cultural origin as a retrieval heuristic. Training on a large Chinese dataset (Chinese-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters. However, this is a shallow fix that fails to generalize to other non-Latin scripts. In contrast, the text-aligned model (XLM-R-L/14) performs well across most languages, with a notable exception for queries in Yoruba (Nigeria). This challenge with low-resource languages persists even in more advanced architectures. For instance, MLLM-based models (Jina-E-v4) employ a LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages. For the vision component, a Vision-Language Model (VLM) is used as the image encoder to improve contextual awareness. However, performance drops for low-resource languages.
Figure 8: Association bias evaluation across four models reveals the limitations of monolingual training. The baseline CLIP (a) shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in Chinese-CLIP (b). In contrast, cross-lingual models like XLM-R (c) and particularly Jina-E-v4 (d) prove far more effective at mitigating this bias and maintaining high semantic relevance across diverse countries.
Figure 9: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) Chinese-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
This behavior is clearly visualized in the UMAP [24] projections of the text embeddings as shown in Figure 9. The baseline CLIP-L/14 model exhibits a fractured embedding space, with non-Latin languages forming distinct clusters far from the main Latin-script cluster. This demonstrates a lack of shared semantic understanding. In the Chinese-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated. In contrast, the explicit alignment model, XLM-R-L/14, successfully unifies the embedding space into a single, language-agnostic cluster, demonstrating a truly shared semantic representation across scripts. The only notable outlier is Yoruba, which was not part of this specific model’s alignment training. The MLLM model, Jina-E-v4, exhibits a similar but distinct pattern: it also forms a single, unified cluster, but the embeddings are more widely dispersed. This suggests a more flexible alignment that may capture finer semantic nuances between languages.
To validate these visual findings numerically, we calculated the silhouette score [25] for each language’s text embeddings. This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix I. This quantitatively reinforces that poor semantic understanding in the text encoder (as visualized by the disparate UMAP clusters) is a key driver of higher cultural association bias.
Both modern MLLM-based models and explicit alignment models drastically reduce the association bias compared to the baselines, achieving SP scores below 0.16. However, neither paradigm consistently outperforms the other on this specific task.
6 Discussion
Our evaluation framework distinguishes between two perspectival biases: prevalence bias, driven by data imbalance, and association bias, arising from learned cultural correlations. Our findings show these are distinct challenges.
Explicit cross-lingual alignment, used by the XLM-R models, is a highly effective strategy, achieving the lowest scores for both prevalence bias (DLBKL) and association bias (SP) by directly enforcing a shared semantic space. While modern MLLMs like Jina-E-v4 also perform well against association bias, the persistence of these issues across all models points to a deeper, unresolved problem: the entanglement of semantic concepts with linguistic and cultural artifacts in the model’s embedding space.
The path forward, therefore, requires a fundamental shift in training strategy. Future work must prioritize training objectives that actively enforce non-association by creating a truly global semantic space. This means designing models to map a semantic query, regardless of its language or cultural origin, to all conceptually relevant images, irrespective of their geographical context. For example, the new method must include a curation process to avoid cross-cultural false negative images being pushed away from their corresponding queries, and utilize data augmentation to help ensure language-agnostic property.
7 Conclusion
In this work, we introduce a framework that distinguishes between two forms of perspectival bias in multimodal retrieval, prevalence bias and association bias, reflecting distinct ways in which a model’s prior shapes its behavior. This conceptual framing is operationalized through our proposed metrics and datasets: DLBKL, which measures rank-aware language prevalence bias, and XCM, which quantifies association bias through cross-cultural image retrieval. Together, these tools enable systematic evaluation of how multimodal large language models inherit and express perspectival biases across languages and cultures.
In image-to-text retrieval, prevalence bias arises when a model favors texts from high-resource languages. This problem is addressed by anchoring other languages to the prevalent ones through cross-lingual alignment. In text-to-image retrieval, association bias arises when a model favors images that are culturally associated with the query language rather than semantically faithful to their content. Such bias cannot be resolved through traditional cross-lingual alignment or by merely exposing the model to a wider range of cultural content during training. Ultimately, our findings call for a more principled approach: one that directly mitigates localized spurious association as a core design principle for models that are not only multilingual but also perform consistently across languages and cultures.
8 Limitations
Our work has several limitations. First, our DLBKL metric measures fairness via distributional parity, not semantic correctness. It therefore cannot distinguish between retrieving irrelevant documents and over-representing a language with relevant ones. Second, the XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures. The benchmark’s coarse-grained semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity.
Acknowledgments
This research has received funding support from the National Science, Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B13F680099] and Chulalongkorn University Artificial Intelligence Institute.
Appendix A Culturally Relevant Concept Identification
To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection. The prompt used in this process is shown in Figure 11 to identify all labels associated with an image. Figure 10.
list 100 concepts that unique and vary in these country including China, India, Japan, saudi arabia, France, German, Brazil,
Kenya, Thailand, USA
like this
{"food":{"China":"Mala Xiang Guo", ..., "Thailand":"Padthai", "USA":"hamburger"}, "costume":{"China":"Hanfu", ...,
"Thailand":"Sabai", "USA":"cowboy"}}
Figure 10: The prompt given to Gemini to generate unique country-specific image concepts.
Appendix B Multi-Concept Detection in Cultural Images
To establish the self-preference cultural bias, the culturally relevant and non-relevant candidate images must not share concepts with the text label. We utilized Gemini2 with the following prompt in Figure 11 to identify all labels associated with an image.
Please classify the following image by assigning them to one or more of the following cultural
categories:
   {category}.
**Comprehensive Output Format (JSON):**
output as JSON for example:
  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
in the categories please order by priority (high to low).
Figure 11: The prompt given to Gemini for multi-label image classification, where the {category} placeholder is dynamically populated with the full list of categories.
Appendix C Language Resources
To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [26] as an approximation. Table 3 and 4 presents the resulting language composition for RQ1 and RQ2, consequently.
Language Type Language ID Distribution (%)
High English en 43.9499
Russian ru 5.7614
German de 5.5691
Medium Japanese ja 4.9152
Chinese-Simpl. zh 4.8778
Spanish es 4.5422
French fr 4.3271
Italian it 2.4060
Portuguese pt 2.3369
Polish pl 1.8744
Dutch nl 1.8083
Indonesian id 1.1759
Turkish tr 1.1274
Czech cs 1.0479
Vietnamese vi 1.0213
Low Korean ko 0.7865
Farsi fa 0.7087
Swedish sv 0.6736
Arabic ar 0.6722
Romanian ro 0.6374
Ukrainian uk 0.6079
Greek el 0.5651
Hungarian hu 0.5082
Danish da 0.4792
Thai th 0.4269
Finnish fi 0.3649
Norwegian no 0.3135
Hebrew he 0.2654
Croatian hr 0.2339
Hindi hi 0.2004
Bengali bn 0.1064
Telugu te 0.0213
Swahili sw 0.0102
Filipino fil 0.0084
Maori mi 0.0014
Cusco Quechua quz 0.0005
Table 3: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ1
Name Full Name Language Resources
USA United States of America English 43.950
UK United Kingdom
AUS Australia
GER Germany German 5.569
CHN China Chinese 4.878
JPN Japan Japanese 4.915
ESP Spain Spanish 4.542
ARG Argentina
FRA France French 4.327
PRT Portugal Portuguese 2.337
BRA Brazil
SAU Saudi Arabia Arabic 0.672
THA Thailand Thai 0.427
IND India Hindi 0.200
KEN Kenya Swahili 0.010
NGA Nigeria Yoruba 0.001
Table 4: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ2
Appendix D Example of Result from Image to Text Retrieval
To further elaborate the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 12. We also provide result of LBKL and DLBKL score at other rank in Table 5.
Figure 12: An Example of Result from Image to Text Retrieval
Model @5 @25 @50 @99
[…]
Vision-Language Contrastive Models
CLIP-L/14 6.904 6.911 4.171 4.182 3.245 3.249 2.658 2.652
Chinese-CLIP-L/14 6.220 6.229 3.793 3.798 3.172 3.168 2.582 2.570
Cross-lingual Alignment Models
XLM-R-L/14 0.939 0.960 0.240 0.246 0.221 0.223 0.214 0.213
XLM-R-B/16plus 0.692 0.713 0.043 0.049 0.030 0.033 0.019 0.019
MLLM-Based Retrieval Embedders
ColQwen2.5-3B-M 2.138 2.164 0.192 0.209 0.114 0.122 0.088 0.089
ColQwen2.5-7B-M 2.373 2.397 0.212 0.232 0.127 0.138 0.083 0.089
ColQwen2.5-v0.2 5.958 5.974 1.375 1.424 0.633 0.676 0.377 0.410
GME-Qwen2-2B 6.014 6.037 0.739 0.804 0.254 0.306 0.140 0.175
GME-Qwen2-7B 3.843 3.875 0.221 0.264 0.094 0.124 0.058 0.077
Jina-E-v4 2.859 2.886 0.157 0.186 0.072 0.093 0.045 0.059
Table 5: Image-to-text retrieval bias on Crossmodal-3600, measured by LBKL and DLBKL at various retrieval depths (k).
Appendix E Language and Rank Frequency Diagram
To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix C, showing both their overall retrieval frequency as shown in Figure 13 and their frequency distribution across ranks as shown in Figure 14.
Figure 13: A language frequency histogram of each language group for all model
Figure 14: A frequency of language group at each rank for all model
Appendix F Annotator Guideline
The guideline we provide to the annotators is to remove duplicates across multiple views. If an image depicts the same scene or object with no meaningful change, keep only one copy. Keep images if there is a significant variation. Allowed differences include time of day (e.g., day vs. night), viewpoint or angle (if the perspective changes enough that visual elements in the image are noticeably different). Minor or trivial variations are not allowed as they would be too similar. This 397 includes slight shifts, crops, or zooms of the same scene.
Appendix G UMAP Analysis for Self-Preference Cultural Bias
To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 15 and 16. The text embeddings cluster strongly by language, a proximity that supersedes semantic content. Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias. While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models. The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster. This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
Figure 15: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CLIP-L/14 model applied to our dataset.
Figure 16: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the Chinese-CLIP-L/14 model applied to our dataset.
Appendix H Full Official Model Name
In this paper, we use aliases for the model names for conciseness; the full names are provided in Table 6.
Alias Used in Paper Full Model Name Parameter
CLIP-L/14 clip-vit-large-patch141 427.6M
Chinese-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
ColQwen2.5-v0.2 ColQwen2.5-v0.23 3814.8M
ColQwen2.5-3B-M ColQwen2.5-3b-multilingual-v1.03 3994.6M
ColQwen2.5-7B-M ColQwen2.5-7b-multilingual-v1.03 8071.1M
GME-Qwen2-2B gme-Qwen2-VL-2B-Instruct4 2209.0M
GME-Qwen2-7B gme-Qwen2-VL-7B-Instruct4 7070.6M
Jina-E-v4 jina-embeddings-v45 3934.7M
XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
The models are based on the following works: 1) Radford et al. [2] for CLIP-L/14; 2) Yang et al. [27] for Chinese-CLIP-L/14; 3) Faysse et al. [14] for ColQwen2 models; 4) Zhang et al. [15] for GME-Qwen2 models; 5) Günther et al. [20] for Jina-E-v4; and 6) Carlsson et al. [10] for XLM-R-VL models.
Table 6: Aliases Used in Paper and Corresponding Full Model Names and Parameters
Appendix I Correlation Analysis for Self-Preference Cultural Bias
We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix G, impacts cross-modal retrieval. To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with self-preference cultural bias score (SP) cultural bias as shown in Table 8. We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores. For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 17. Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder as shown in Table 7.
Figure 17: Comparison of Self-Preference Cultural Bias with Text and Image Silhouette
Model TSC ISC
CLIP-L/14 0.83 0.55
Chinese-CLIP-L/14 -0.26 0.58
XLM-R-VL-B/16 0.98 -0.27
XLM-R-VL-L/14 0.94 0.05
Jina-E-v4 0.86 0.16
GME-Qwen2-2B 0.80 0.09
GME-Qwen2-7B 0.64 0.07
Average 0.68 0.18
Table 7: This table presents a Pearson correlation analysis between model performance and bias. We measure the correlation between association and the quality of data clusters (via Silhouette Score) in both the text embedding space (TSC) and the image embedding space (ISC).
Model Metrics Country
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
CLIP-L/14 SP $\downarrow$ 0.01 0.02 0.02 0.76 2.63 1.94 0.24 0.19 0.34 0.39 0.51 10.71 8.09 15.88 2.04 2.27
TS $\downarrow$ 0.05 0.05 0.05 0.02 0.16 -0.05 -0.01 -0.02 -0.02 -0.01 -0.01 0.25 0.23 0.23 0.03 0.13
IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.04 0.01 0.01 0.00 0.00 0.01 0.05 0.03 0.03 0.02 0.03
Chinese-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
TS $\downarrow$ -0.05 -0.05 -0.05 0.00 0.13 -0.07 -0.03 -0.03 -0.03 -0.01 -0.01 -0.40 0.93 -0.37 0.00 0.03
IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.00 0.00 0.00 0.04 0.05 0.04 0.02 0.01
Jina-E-v4 SP $\downarrow$ 0.02 0.03 0.02 0.05 0.04 0.04 0.04 0.03 0.05 0.04 0.04 0.05 0.02 0.12 0.49 0.93
TS $\downarrow$ -0.01 -0.01 -0.01 -0.02 0.01 0.00 -0.01 -0.02 -0.02 -0.02 -0.02 0.02 0.01 0.00 0.00 0.13
IS $\downarrow$ -0.01 0.00 -0.01 -0.01 0.00 0.01 0.00 0.00 -0.01 -0.01 -0.01 0.02 0.00 0.01 0.01 0.00
XLM-R-L/14 SP $\downarrow$ 0.05 0.04 0.07 0.05 0.04 0.07 0.08 0.08 0.09 0.08 0.07 0.05 0.04 0.04 0.11 0.61
TS $\downarrow$ -0.18 -0.18 -0.18 -0.11 -0.07 -0.10 -0.16 -0.16 -0.16 -0.12 -0.12 -0.07 -0.06 -0.17 -0.14 0.44
IS $\downarrow$ 0.01 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.01 0.01 0.01 0.06 0.04 0.03 0.02 0.03
XLM-R-B/16plus SP $\downarrow$ 0.05 0.04 0.06 0.04 0.04 0.05 0.05 0.06 0.07 0.06 0.06 0.04 0.02 0.05 0.10 0.66
TS $\downarrow$ -0.24 -0.24 -0.24 -0.21 -0.19 -0.19 -0.24 -0.23 -0.23 -0.23 -0.23 -0.19 -0.19 -0.23 -0.23 0.49
IS $\downarrow$ 0.01 0.01 0.00 0.00 0.01 0.01 0.00 0.00 -0.01 0.00 -0.01 0.03 0.02 0.03 0.02 0.00
GME-Qwen2-2B SP $\downarrow$ 0.02 0.03 0.02 0.10 0.04 0.04 0.05 0.05 0.08 0.08 0.09 0.14 0.10 0.24 1.24 2.02
TS $\downarrow$ 0.02 0.02 0.02 0.04 0.02 0.00 0.04 0.01 0.01 0.03 0.03 0.01 0.03 0.00 0.03 0.15
IS $\downarrow$ 0.00 0.00 0.01 0.00 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.02 0.01 0.01
GME-Qwen2-7B SP $\downarrow$ 0.03 0.02 0.01 0.14 0.04 0.05 0.07 0.07 0.07 0.09 0.13 0.14 0.08 0.12 0.62 1.88
TS $\downarrow$ 0.04 0.04 0.04 0.05 0.02 0.02 0.07 0.02 0.02 0.04 0.04 0.09 0.11 0.03 0.04 0.14
IS $\downarrow$ 0.00 0.00 0.01 0.00 0.01 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.01 0.01 0.01
Table 8: Cross-Country and Cross-Model Comparison of Language and Cultural Bias Metrics. This table presents the results for the Self-Preference Cultural Bias score (SP), Text Silhouette (TS), and Image Silhouette (IS) scores across various multimodal retrievers for a selection of countries.
Appendix J Dataset Statistics
The distribution of cultural concepts in the XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
Country Number of Samples
Argentina 771
Australia 721
Brazil 724
China 727
France 760
Germany 744
India 774
Japan 944
Kenya 600
Nigeria 773
Portugal 824
Saudi Arabia 619
Spain 841
Thailand 649
UK 644
USA 609
Average 733
Table 9: Dataset Statistics of XCM dataset per Country
Concepts (A-G)
airlines 24
airport 65
alcohol drink 26
ancient city 112
ancient craft 111
ancient painting 139
animal 100
architecture 107
art 89
artwork 26
bag 20
bakery 47
banknotes 69
bathroom 30
bedroom 77
boat 99
bracelet 62
building 196
bus 92
bus station 56
capital 147
celebrity 66
child 69
Concepts (C-M)
cinema 54
coin 182
combination food 72
congress 127
costume 119
craft 98
dance 145
deep fried food 30
department store 21
dessert 66
devil 43
diningroom 49
doll 131
drink 31
dry heat food 21
embroidery style 133
fashion 57
festival 145
fire station 162
folk tale 38
folklore character 90
food 52
football player 104
Concepts (F-M)
formal uniform 117
fountain style 91
funeral 141
game 76
gas station 63
gathering place 92
ghost 19
graduated uniform 78
hat 95
headwear 106
historical event 89
historical figure 81
historical image 120
hot pot concept 66
hotel 70
house 86
instrument 58
lottery tickets 48
mailbox 81
major mountain range 52
major religious site 97
major river 57
map 32
market 74
Concepts (M-P)
marriage ceremony 95
martial art 96
mask 104
military parade 189
moist heat food 34
museum 99
music band 95
mythical creature 56
mythological figure 68
native inhabitants 163
natural landmark 152
necklace 63
night view 110
older 38
painting 128
pants 30
people 136
poaching food 25
police station 158
popular street food 98
pottery style 150
priest 68
prime minister 86
Concepts (R-S)
religious building 123
restaurant 38
ritual 108
rural dwelling 127
sacred object 101
school 120
series 40
shirt 64
shopping mall 91
singer 56
snack 56
social custom 147
soldier 167
sport 79
stageplay 108
statue 106
street entertainment 111
street sign 81
street vendor cart 32
street view 86
symbolic bird 84
symbolic plant 72
Concepts (T-Z)
tattoo style 59
taxi 102
tea culture 85
textile pattern 56
tourist attraction 130
toy 66
train 116
train station 128
tree 94
tv program 43
unique art form 119
unique cuisine trait 43
unique food ingredient 22
unique natural phenomenon 102
unique transportation 75
university 136
wall painting 65
warrior 79
weapon 75
wedding 108
writing character 15
zoo 79
Table 10: Distribution of Concepts and Image Counts
Model Metrics Country
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
CLIP-L/14 $M_{\text{sem}}(\%)\uparrow$ 95.73 94.57 94.73 52.42 25.17 31.07 75.53 78.00 69.65 65.78 61.33 7.75 10.48 5.56 27.83 24.19
$M_{\text{cul}}(\%)\downarrow$ 1.31 2.02 2.22 39.65 66.16 60.34 18.29 15.10 23.61 25.85 31.35 83.04 84.75 88.24 56.67 54.85
$M_{\text{non}}(\%)\downarrow$ 2.96 3.42 3.05 7.93 8.67 8.59 6.18 6.90 6.74 8.37 7.32 9.21 4.78 6.20 15.50 20.96
Chinese-CLIP -L/14 $M_{\text{sem}}(\%)\uparrow$ 94.25 89.44 90.57 40.05 93.40 84.20 65.79 61.95 60.83 56.55 57.04 14.38 21.88 10.47 27.00 25.10
$M_{\text{cul}}(\%)\downarrow$ 3.28 5.75 4.44 44.62 2.48 10.92 21.97 24.73 26.07 31.80 29.56 70.11 55.78 73.00 53.83 56.02
$M_{\text{non}}(\%)\downarrow$ 2.46 4.81 4.99 15.32 4.13 4.88 12.24 13.32 13.10 11.65 13.40 15.51 22.34 16.54 19.17 18.89
Jina-E-v4 $M_{\text{sem}}(\%)\uparrow$ 95.89 95.65 95.98 92.47 94.22 92.79 94.34 93.22 91.70 91.99 94.20 91.60 95.38 86.30 56.50 36.74
$M_{\text{cul}}(\%)\downarrow$ 1.97 2.48 1.66 4.97 3.44 3.92 4.08 3.09 4.80 4.13 3.73 4.68 2.00 9.95 27.83 34.15
$M_{\text{non}}(\%)\downarrow$ 2.13 1.86 2.36 2.55 2.34 3.29 1.58 3.69 3.50 3.88 2.07 3.72 2.62 3.75 15.67 29.11
XLM-R-L/14 $M_{\text{sem}}(\%)\uparrow$ 89.66 91.15 89.46 90.32 92.30 89.40 87.63 85.37 86.12 87.86 87.98 90.95 93.07 89.41 80.33 40.49
$M_{\text{cul}}(\%)\downarrow$ 4.11 4.04 5.83 4.44 4.13 6.26 6.58 6.54 7.39 7.16 5.80 4.85 3.54 3.75 8.50 24.71
$M_{\text{non}}(\%)\downarrow$ 6.24 4.81 4.72 5.24 3.58 4.35 5.79 8.09 6.49 4.98 6.22 4.20 3.39 6.85 11.17 34.80
XLM-R-B /16Plus $M_{\text{sem}}(\%)\uparrow$ 91.95 91.30 90.71 93.95 94.50 91.62 90.66 88.47 87.81 90.05 89.36 92.57 95.22 91.09 83.17 40.88
$M_{\text{cul}}(\%)\downarrow$ 4.11 3.88 5.13 4.03 3.58 4.56 4.34 5.47 5.97 5.70 5.52 3.88 2.00 4.91 8.33 26.78
$M_{\text{non}}(\%)\downarrow$ 3.94 4.81 4.16 2.02 1.93 3.82 5.00 6.06 6.23 4.25 5.11 3.55 2.77 4.01 8.50 32.34
GME-Qwen2 -2B-Instruct $M_{\text{sem}}(\%)\uparrow$ 96.06 95.34 96.95 87.77 94.09 93.43 92.76 90.49 88.59 90.05 88.95 84.49 89.06 76.87 37.17 25.87
$M_{\text{cul}}(\%)\downarrow$ 2.30 2.48 1.66 8.47 3.58 3.92 4.74 4.64 7.26 7.04 7.87 11.63 8.63 18.48 46.00 52.26
$M_{\text{non}}(\%)\downarrow$ 1.64 2.17 1.39 3.76 2.34 2.65 2.50 4.88 4.15 2.91 3.18 3.88 2.31 4.65 16.83 21.86
GME-Qwen2 -7B-Instruct $M_{\text{sem}}(\%)\uparrow$ 95.40 95.34 96.53 84.68 95.05 92.47 90.39 90.73 90.27 88.35 85.50 85.46 90.91 85.92 55.17 29.62
$M_{\text{cul}}(\%)\downarrow$ 2.46 2.33 0.97 11.69 3.44 4.77 6.45 5.95 6.10 7.77 11.46 11.63 7.09 10.34 34.00 55.76
$M_{\text{non}}(\%)\downarrow$ 2.13 2.33 2.50 3.63 1.51 2.76 3.16 3.33 3.63 3.88 3.04 2.91 2.00 3.75 10.83 14.62
Table 11: Cross-Country and Cross-Model Comparison of Win Percentages. This table presents the results for the Semantically Relevant, Culturally Relevant, and Non-Relevant Win Percentages across various multimodal retrievers for a selection of countries.
Appendix K Computational Resource
The experiment is performed with a single A100 GPU for approximately 3 gpu hours for each model or 54 hours in total with library version of colpali-engine 0.3.13.dev1+g9bee9b2b7, transformers 4.53.3 for most experiment except, GME models are inferenced under transformers 4.51.3
Appendix L Authoring and Implementation Tools
In preparing this manuscript, we utilized several generative large language models. For language editing and stylistic refinement, we employed Google’s Gemini 2.5-flash, along with models from xAI’s Grok family (e.g., Grok-3 Expert and Fast variants). For assistance with code implementation, scripting, and debugging, we used a model from Anthropic’s Claude series (e.g., Claude 4.0 Sonnet).
Appendix M Detailed Results
The full details of RQ2 experiment including all win rate of all models are illustrated in the Table 11.
Appendix N 3XCM Dataset Benckmark Samples
This research provides an association evaluation benchmark and image metadata. Examples of the benchmark and image metadata are shown in Figure 18 and Figure 19, respectively.
Figure 18: Examples of 3XCM dataset benchmark
Figure 19: Metadata for image of 3XCM dataset benchmark